Use the data in the following​ table, which lists​ drive-thru order accuracy at popular fast food chains. Assume that orders are randomly selected from those included in the table. ​Drive-thru Restaurant A B C D Order Accurate 323 274 244 130 Order Not Accurate 32 56 36 19 If two orders are​ selected, find the probability that they are both from Restaurant D. a. Assume that the selections are made with replacement. Are the events​ independent? b. Assume that the selections are made without replacement. Are the events​ independent?
Answer #1

Orders from D = 130 + 19 = 149

Total orders = 323 + 274 + 244 + 130 + 32 + 56 + 36 + 19 = 1114

Hence,

a) With replacement:

P(Both from D) = (149/1114) * (149/1114) = 0.0180

Yes, Independent

b) Without replacement:

P(Both from D) = (149/1114)*(148/1113) = 0.0178

No, not independent